# Egorov-Type Semiclassical Limits for Open Quantum Systems with a Bi-Lindblad Structure

### v1

@@ 3.4 Example: a Poisson–Nijenhuis bi-Hamiltonian system with two nontrivial contact integrals
with coordinates $(x^{1},x^{2},p_{1},p_{2})$ and canonical one-form $\theta=p_{i}\,\mathrm{d}x^{i}$. Throughout we use the convention $\omega=-\mathrm{d}\theta$. The associated Poisson tensor is
@@
The Liouville vector field
$\Delta=p_{i}\frac{\partial}{\partial p_{i}}$
satisfies $\iota_{\Delta}\omega=-\theta$, and $\lambda_{1},\lambda_{2}$ are both $1$-homogeneous:
@@
Then $\mathcal{L}_{\Delta}H=H$ and, since $\{\lambda_{1},\lambda_{2}\}_{\Lambda}=0$, each $\lambda_{i}$ is a first integral of $X_{H}$. Thus $(M,\theta,H)$ is a $1$-homogeneous integrable system in the sense of Definition 2.3.
@@
Since $\Delta=r\partial_{r}$ is transverse to the level sets of $r$, the hypersurface $C$ is transverse to the Liouville flow. The restriction
@@
is a contact form, since
@@
on a dense open subset (in fact everywhere). Therefore $(C,\alpha,h)$ is a completely integrable contact system in the sense of Definition 2.4.
@@ 4 Lindblad dynamics compatible with contact geometry
We now move to the quantum setting and introduce a class of Lindblad generators that are compatible with the contact integrable systems constructed in Section 3. For definiteness, we work with finite-dimensional Hilbert spaces; in this context, $\mathcal{B}(\mathcal{H})$ denotes the full $C^{\ast}$-algebra of bounded operators on $\mathcal{H}$ (which coincides with the space of all linear operators), all operator topologies coincide, and domain issues do not arise. Most of the algebraic definitions and arguments below extend to separable infinite-dimensional Hilbert spaces under standard technical assumptions on domains and continuity of the corresponding quantum dynamical semigroups.
@@
In the finite-dimensional setting, a state is described by a density matrix $\rho$. Fixing an orthonormal basis $\{|j\rangle\}_{j=1}^{d}$ (often selected by the measurement scheme or by the spectral decomposition of a reference Hamiltonian), we write $\rho_{jk}=\langle j|\rho|k\rangle$. The coherences are the off-diagonal entries $\rho_{jk}$ with $j\neq k$, which encode relative phases and interference effects, while the diagonal entries $\rho_{jj}$ represent populations, i.e., classical probabilities of outcomes in that basis (in the standard terminology of open quantum systems, see e.g., [BreuerPetruccione]). A dynamics exhibits decoherence (in that basis) when the off-diagonal entries are damped in time, typically driving $\rho(t)$ towards a state that is approximately diagonal in the chosen basis. In open quantum systems, decoherence is not an ad hoc ingredient but a robust consequence of the coupling to an environment: while the total evolution on system-plus-environment may be unitary, the reduced dynamics on the system alone is generally non-unitary and may suppress coherences.
There are two equivalent ways to represent the same quantum evolution. It is convenient to introduce a unified notation for expectation values. For an observable $A$ at time $t$, we define
$\mathrm{E}(A)(t)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Tr}\big(\rho(t)A\big)$
in the Schrödinger picture, or equivalently
$\mathrm{E}(A)(t)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\operatorname{Tr}\big(\rho_{0}A(t)\big)$
in the Heisenberg picture.
In the Schrödinger picture, the state evolves according to
@@
on the commutative subalgebra generated by the integrals. This viewpoint allows one to identify conserved quantities and invariant subalgebras directly at the level of the observable algebra $\mathcal{B}(\mathcal{H})$, which is particularly well suited to the geometric framework developed here.
By contrast, the Schrödinger picture is often preferable when addressing state-preparation problems, entropy production, contractivity properties of the quantum dynamical semigroup, or the detailed long-time behavior of the density operator $\rho(t)$; see, for instance, [Lindblad1976, BreuerPetruccione] and the references therein. Both pictures are, of course, mathematically equivalent, but emphasize different structural aspects of open quantum dynamics.
@@ 4.1 GKSL generators and Heisenberg picture
It is well known that such $\mathcal{L}$ are precisely the generators of one-parameter semigroups $\Phi_{t}=e^{t\mathcal{L}},\,t\geq 0$, of completely positive and trace-preserving (CPTP) maps on $\mathcal{S}(\mathcal{H})$. Conversely, any norm-continuous CPTP semigroup on $\mathcal{S}(\mathcal{H})$ has a generator of the form (4.1) (see [BreuerPetruccione], [Lindblad1976] for instance).
@@
Since the trace pairing $\langle\rho,A\rangle=\mathrm{Tr}(\rho A)$ is nondegenerate, $\mathcal{L}^{\dagger}$ is uniquely determined by this relation.
@@
and, by the defining relation of $\mathcal{L}^{\dagger}$ and the nondegeneracy of the trace pairing, this shows (4.2). ∎
@@
Since the trace product is non-degenerate, the right-hand side vanishes for every $\rho$ if and only if $\mathcal{L}^{\dagger}(O)$. ∎
@@
For any Lindblad operator, every matrix proportional to the identity is a constant of the motion, since $\mathbb{I}\in\ker\mathcal{L}^{\dagger}$. Hence, constants of the motion always form an unital subalgebra of $\mathcal{B}(\mathcal{H})$, of dimension at least $1$.
@@
since $L_{k},O]=[L_{k}^{\dagger},O]=0$ implies $O$ commutes with both $L_{k}$ and $L_{k}^{\dagger}$, and hence with $L_{k}^{\dagger}L_{k}$. Similarly,
@@
This will be our basic mechanism to quantize the classical dissipated quantities of a contact integrable system: we will construct Lindblad generators whose Hamiltonian and dissipative parts lie in the commutant of the algebra generated by the quantum integrals. $\diamond$
@@
This dynamics is known as pure dephasing: it suppresses the off-diagonal matrix elements of $\rho$ in the eigenbasis of $\sigma_{z}$ (i.e. quantum coherences) while leaving the diagonal entries (the populations) unchanged. Although dissipation is present, no energy exchange with the environment occurs.
@@
Finally, recall that the identity operator $\mathbb{I}$ always satisfies $\mathcal{L}^{\dagger}(\mathbb{I})=0$, expressing trace preservation of the Lindblad dynamics. Therefore, this system has a nontrivial subalgebra $\operatorname{span}\{\sigma_{z},\mathbb{I}\}\subset\mathcal{B}(\mathcal{H})$ of conserved observables, despite the presence of irreversible dissipation.
@@ 4.2 Invariant commutative subalgebras and constants of motion
From a geometric viewpoint, this selection mechanism parallels the emergence of dissipated quantities in contact Hamiltonian systems: although the full dynamics is non-Hamiltonian and generically dissipative, certain observables retain a privileged status encoded by algebraic invariance conditions. This analogy will be made precise in the following sections by requiring that the invariant quantum observables coincide, in the semiclassical limit, with the dissipated quantities of a completely integrable contact system. $\diamond$
@@
for some $C>0$. This exponential bound is the precise sense in which the dissipative dynamics suppresses the noncommutative (coherence) sector while leaving $\mathcal{A}$ pointwise invariant.
@@
More generally, this example illustrates a robust structural phenomenon: for suitable GKSL generators, there exist nontrivial commutative $C^{\ast}$-subalgebras $\mathcal{A}\subset\mathcal{B}(\mathcal{H})$, with $\mathcal{A}\neq\operatorname{span}(\mathbb{I})$, that are invariant under the Heisenberg evolution and lie entirely in $\ker\mathcal{L}^{\dagger}$.
@@ 4.3 Contact-compatible Lindblad generators
Consider the algebra $(\mathscr{C}^{\infty}(C),\{\cdot,\cdot\}_{\alpha})$, where $\{\cdot,\cdot\}_{\alpha}$ denotes de Jacobi bracket defined by $\alpha$. Let $\mathcal{A}_{\mathrm{cl}}\subset\mathscr{C}^{\infty}(C)$ denote the commutative subalgebra generated by $\{I_{0},\dots,I_{n}\}$. We refer to $\mathcal{A}_{\mathrm{cl}}$ as the Jacobi commutative algebra of the integrable system. This terminology reflects the fact that $\mathcal{A}_{\mathrm{cl}}$ plays, for contact dynamics, the same role as the Poisson-commutative algebra of integrals in Liouville-integrable Hamiltonian systems: it encodes the distinguished observables whose mutual relations are governed by the underlying Jacobi structure.
@@
A contact system is naturally encoded by its symplectization. Given a contact manifold $(C,\alpha)$, its symplectization is
$(M,\omega)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=(C\times\mathbb{R}_{>0},\,\omega\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=d(r\alpha)),$
with Liouville (Euler) vector field $\Delta=r\partial_{r}$. Any function $f\in\mathscr{C}^{\infty}(C)$ admits a canonical 11–homogeneous lift on $M$, for instance
$f^{\Sigma}(x,r)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=r\,f(x),$
and the Jacobi bracket on $(C,\alpha)$ is recovered from the Poisson bracket on $(M,\omega)$ by restriction to the hypersurface $\{r=1\}$:
$\{f,g\}_{\alpha}\;=\;\{f^{\Sigma},g^{\Sigma}\}_{\omega}\big|_{r=1}.$
In particular, the contact Hamiltonian dynamics on $C$ is the reduction (or, equivalently, the restriction to a transverse section $r=\mathrm{const}$) of the homogeneous Hamiltonian dynamics on $M$.
@@
For this reason, the correct correspondence between classical and quantum structures should be understood as follows: the quantization map $Q_{\hbar}$ assigns operators to the nonconstant generators of the Jacobi-commutative algebra $\mathcal{A}_{\mathrm{cl}}$, while the unit is added automatically on the quantum side. The nontrivial content of contact compatibility and of quantum constants of motion therefore lies entirely in the quantization of the dissipated quantities, not in the unit element.
@@
Let $\mathcal{A}_{\mathrm{cl}}=\mathrm{span}\{I\}$ be the Jacobi-commutative subalgebra generated by a single dissipated quantity $I$. Any semiclassical quantization $Q_{\hbar}$ maps the nontrivial generator $I$ to a self-adjoint operator $\widehat{I}_{\hbar}$. The corresponding quantum algebra is then taken to be the unital $C^{\ast}$-algebra
@@
A contact-compatible Lindblad generator is precisely one whose Heisenberg adjoint leaves $\mathcal{A}_{\hbar}$ pointwise invariant and satisfies $\mathcal{L}^{\dagger}(\widehat{I}_{\hbar})=0$. In this minimal setting, contact compatibility reduces to the existence of a nontrivial invariant commutative subalgebra generated by a single dissipated quantity beyond the scalars. This example highlights a general structural feature: while the constant function $1$ plays a nontrivial role in the Jacobi algebra at the classical level, its quantum counterpart is automatically invariant. The content of contact compatibility therefore lies entirely in the quantization of the nonconstant dissipated quantities.
@@
$\{I_{i},I_{j}\}_{\alpha}=0\quad\text{for all }i,j,\qquad\{h,I_{i}\}_{\alpha}=-\lambda_{i}I_{i}$
for some contact Hamiltonian $h$ and real constants $\lambda_{i}$.
@@
Assumption (2) ensures that the Hamiltonian and the Lindblad operators are chosen inside the commutant of the $C^{\ast}$-algebra generated by the $\widehat{I}_{i,\hbar}$. Therefore, the operators $\widehat{I}_{i,\hbar}$ generate a commutative $C^{\ast}$-subalgebra $\mathcal{A}_{\hbar}\subset\mathcal{B}(\mathcal{H}_{\hbar})$, which is unital by construction, since any $C^{\ast}$-subalgebra of $\mathcal{B}(\mathcal{H}_{\hbar})$ contains the identity operator.
@@ 4.4 Bi-Lindblad structures
Let $(C,\alpha,h;I_{0},\dots,I_{n})$ and $(\mathcal{H}_{\hbar},Q_{\hbar})$ be as in Definition 4.14. Assume in addition that the contact system arises from a bi-Hamiltonian Poisson–Lie system $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ as in Section 3, so that the classical vector field $X$ is Hamiltonian both for $\pi_{0}$ and $\pi_{1}$, and the hierarchy of integrals $\{I_{j}\}$ ultimately comes from the common bi-Hamiltonian invariants on $G$.
@@
Heuristically, (BL1) mirrors the classical requirement that any linear combination $\pi_{\lambda}=\pi_{1}-\lambda\pi_{0}$ of compatible Poisson bivectors is still a Poisson tensor: here convex combinations of the two generators remain physically admissible Lindblad generators. Condition (BL2) captures the fact that the same family of classical integrals $I_{j}$ arises from the bi-Hamiltonian hierarchy and should therefore be conserved by both quantum evolutions. Finally, (BL3) encodes the requirement that, in the semiclassical limit, each branch $\mathcal{L}^{(a)}_{\hbar}$ recovers the contact dynamics associated with a different classical Hamiltonian $H_{a}$, in analogy with the way a bi-Hamiltonian vector field is Hamiltonian for both $\pi_{0}$ and $\pi_{1}$.
@@
Verification of (CC1). Since the Heisenberg adjoint is affine,
@@
We illustrate Definition 4.21 on the two-level system $\mathcal{H}=\mathbb{C}^{2}$, where the invariant commutative sector is generated by $\sigma_{z}$. Fix parameters $\omega_{0},\omega_{1}\in\mathbb{R}$ and $\gamma>0$, and define two GKSL generators on $\mathcal{B}(\mathbb{C}^{2})$ (not forming a semiclassical family in $\hbar$, but sufficient to illustrate the algebraic content of Definition 4.21):
@@
Note that $\mathcal{L}^{(0)}$ describes a purely unitary evolution, while $\mathcal{L}^{(1)}$ corresponds to a unitary evolution plus pure dephasing in the $\sigma_{z}$ basis. For every $\lambda\in[0,1]$, let
@@
Thus a bi-Lindblad structure provides, for each $\lambda\in[0,1]$, a specific choice of contact-compatible Lindblad generator whose semiclassical limit reflects a different point on the classical bi-Hamiltonian pencil. In the deformed Euler top example of Section 5, we shall see explicitly how the quantum generators $\mathcal{L}^{(0)}_{\hbar}$ and $\mathcal{L}^{(1)}_{\hbar}$ can be chosen so that their common quantum integrals $\widehat{I}_{j,\hbar}$ quantize the contact dissipated quantities $I_{j}$ obtained in Section 3, thereby realizing a concrete bi-Lindblad structure in the above sense.
@@
Before stating the Egorov-type criterion, we briefly recall what we mean by semiclassical Weyl quantization in the standard pseudodifferential setting. Assume that the exact symplectic manifold $(M,\omega)$ is globally symplectomorphic to $\mathbb{R}^{2d}$ (with $d=n+1$) with canonical coordinates $(x,\xi)\in\mathbb{R}^{\mathrm{d}}\times\mathbb{R}^{d}$ and symplectic form ${\omega=\sum_{j=1}^{d}\mathrm{d}\xi_{j}\wedge dx_{j}}$.
@@
in operator norm for $a,b\in S^{0}(M)$ under standard symbol assumptions. This is the pseudodifferential backbone behind the Dirac condition and the Egorov-type limit used below. We refer to [Zworski2012, DimassiSjoestrand1999] for comprehensive accounts of semiclassical Weyl quantization, symbol classes $S^{m}$, the Moyal product, and the Egorov theorem. See also [Folland1989, Martinez2002] for the Weyl calculus and its microlocal properties. From the perspective of deformation quantization, the Moyal product and the Dirac condition are discussed for instance in [Gutt2000].
@@
$(C,\alpha,h)=(\mathbb{R}^{3}(q,p,z),\,\mathrm{d}z-p\,\mathrm{d}q,\,h=z-p),\qquad I_{1}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=e^{-z}.$
@@
$\mathcal{A}_{\mathrm{cl}}=\mathrm{Jacobi\text{-}alg}\langle I_{0}=h,\,I_{1}=e^{-z}\rangle\subset\mathscr{C}^{\infty}(C).$
@@
As a consequence, the Lindblad operators belong to $\mathcal{A}_{\hbar}$ and hence
@@
$\widehat{I}_{0,\hbar}=\widehat{h}_{\hbar},\qquad\widehat{I}_{1,\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(I_{1}^{\Sigma}),\qquad I_{1}=e^{-z}.$

### v2

@@ 3.4 Example: a Poisson–Nijenhuis bi-Hamiltonian system with two non-trivial contact integrals
with coordinates $(x^{1},x^{2},p_{1},p_{2})$ equipped with the canonical one-form $\theta=p_{i}\,\mathrm{d}x^{i}$ and the canonical symplectic form $\omega=-\mathrm{d}\theta=\mathrm{d}x^{i}\wedge\mathrm{d}p_{i}$. The associated Poisson tensor is
@@
The Liouville vector field reads
$\Delta=p_{i}\frac{\partial}{\partial p_{i}}\,$,
and $\lambda_{1},\lambda_{2}$ are both $1$-homogeneous:
@@
such that each $\lambda_{i}$ is a first integral of $X_{H}$. Thus $(M,\theta,H)$ is a $1$-homogeneous integrable system in the sense of Definition 2.3.
@@
Since $\Delta=r\partial_{r}$ is transverse to the level sets of $r$, the hypersurface $C$ is transverse to the Liouville flow. By Lemma 3.5, the restriction
@@
is a contact form. Indeed,
@@
on a dense open subset (in fact everywhere). Therefore $(C,\alpha,h)$ is a completely integrable contact system in the sense of Definition 2.4. As a matter of fact, it is the system from Example 2.10.
@@ 4 Lindblad dynamics compatible with contact geometry
We now move to the quantum setting and introduce a class of Lindblad generators that are compatible with the contact integrable systems constructed in Section 3. For simplicity’s sake, we work with finite-dimensional Hilbert spaces; in this context, $\mathcal{B}(\mathcal{H})$ denotes the full $C^{\ast}$-algebra of bounded operators on $\mathcal{H}$ (which coincides with the space of all linear operators), all operator topologies coincide, and domain issues do not arise. Most of the algebraic definitions and arguments below extend to separable infinite-dimensional Hilbert spaces under standard technical assumptions on domains and continuity of the corresponding quantum dynamical semigroups.
@@
In the finite-dimensional setting, a state is described by a density matrix $\rho$. Fixing an orthonormal basis $\{|j\rangle\}_{j=1}^{d}$ (often selected by the measurement scheme or by the spectral decomposition of a reference Hamiltonian), we write $\rho_{jk}=\langle j|\rho|k\rangle$. The coherences are the off-diagonal entries $\rho_{jk}$ with $j\neq k$, which encode relative phases and interference effects, while the diagonal entries $\rho_{jj}$ represent populations, i.e., classical probabilities of outcomes in that basis (in the standard terminology of open quantum systems, see e.g., [6]). A dynamics exhibits decoherence (in that basis) when the off-diagonal entries are damped in time, typically driving $\rho(t)$ towards a state that is approximately diagonal in the chosen basis. In open quantum systems, decoherence is not an ad hoc ingredient but a robust consequence of the coupling to an environment: while the total evolution on system-plus-environment may be unitary, the reduced dynamics on the system alone is generally non-unitary and may suppress coherences.
There are two equivalent ways to represent the same quantum evolution. In the Schrödinger picture, the state evolves according to
@@
on the commutative subalgebra generated by the functions in involution. This viewpoint allows one to identify conserved quantities and invariant subalgebras directly at the level of the observable algebra $\mathcal{B}(\mathcal{H})$, which is particularly well suited to the geometric framework developed here.
By contrast, the Schrödinger picture is often preferable when addressing state-preparation problems, entropy production, contractivity properties of the quantum dynamical semigroup, or the detailed long-time behavior of the density operator $\rho(t)$; see, for instance, [42, 6] and the references therein. Both pictures are, of course, mathematically equivalent, but emphasize different structural aspects of open quantum dynamics.
@@ 4.1 GKSL generators and Heisenberg picture
It is well known that such $\mathcal{L}$ are precisely the generators of one-parameter semigroups $\Phi_{t}=e^{t\mathcal{L}},\,t\geq 0$, of completely positive and trace-preserving (CPTP) maps on $\mathcal{S}(\mathcal{H})$. Conversely, any norm-continuous CPTP semigroup on $\mathcal{S}(\mathcal{H})$ has a generator of the form (4.1) (see [6, 42] for instance).
@@
Since the trace pairing $\langle\rho,A\rangle=\mathrm{Tr}(\rho A)$ is non-degenerate, $\mathcal{L}^{\dagger}$ is uniquely determined by this relation.
@@
and, by the defining relation of $\mathcal{L}^{\dagger}$ and the non-degeneracy of the trace pairing, this shows (4.2). ∎
@@
Since the trace product is non-degenerate, the right-hand side vanishes for every $\rho$ if and only if $\mathcal{L}^{\dagger}(O)=0$. ∎
@@
For any Lindblad operator, every matrix proportional to the identity is a constant of the motion, since $\mathbb{I}\in\ker\mathcal{L}^{\dagger}$ for any $\mathcal{L}^{\dagger}$ of the form (4.2). Hence, constants of the motion always form an unital subalgebra of $\mathcal{B}(\mathcal{H})$, of dimension at least $1$.
@@
since $L_{k},O]=[L_{k}^{\dagger},O]=0$ implies that $O$ commutes with $L_{k}^{\dagger}L_{k}$. Similarly,
@@
This will be our basic mechanism to quantize the classical dissipated quantities of a contact integrable system: we will construct Lindblad generators whose Hamiltonian and dissipative parts lie in the commutant of the algebra generated by the quantum constants of the motion. $\diamond$
@@
This dynamics is known as pure dephasing: it suppresses the off-diagonal matrix elements of $\rho$ in the eigenbasis of $\sigma_{z}$ (i.e. quantum coherences) while leaving the diagonal entries (the populations) unchanged. Although dissipation is present, no energy exchange with the environment occurs, i.e., $H$ is a constant of the motion.
@@
Finally, recall that the identity operator $\mathbb{I}$ always satisfies $\mathcal{L}^{\dagger}(\mathbb{I})=0$, expressing trace preservation of the Lindblad dynamics. Therefore, this system has a non-trivial subalgebra $\operatorname{span}\{\sigma_{z},\mathbb{I}\}\subset\mathcal{B}(\mathcal{H})$ of conserved observables, despite the presence of irreversible dissipation.
@@ 4.2 Invariant commutative subalgebras and constants of motion
From a geometric viewpoint, this selection mechanism parallels the emergence of dissipated quantities in contact Hamiltonian systems: although the full dynamics is non-Hamiltonian and generically dissipative, certain observables retain a privileged status encoded by algebraic invariance conditions. This analogy will be made precise in the following sections by requiring that the invariant quantum observables coincide, in the semiclassical limit, with the dissipated quantities of a completely integrable contact system.
@@
for some $C>0$. This exponential bound is the precise sense in which the dissipative dynamics suppresses the non-commutative (coherence) sector while leaving $\mathcal{A}$ pointwise invariant.
@@
More generally, this example illustrates a robust structural phenomenon: for suitable GKSL generators, there exist non-trivial commutative $C^{\ast}$-subalgebras $\mathcal{A}\subset\mathcal{B}(\mathcal{H})$, with $\mathcal{A}\neq\operatorname{span}(\mathbb{I})$, that are invariant under the Heisenberg evolution and lie entirely in $\ker\mathcal{L}^{\dagger}$.
@@ 4.3 Contact-compatible Lindblad generators
Consider the algebra $(\mathscr{C}^{\infty}(C),\{\cdot,\cdot\}_{\alpha})$, where $\{\cdot,\cdot\}_{\alpha}$ denotes the Jacobi bracket defined by $\alpha$. Let $\mathcal{A}_{\mathrm{cl}}\subset\mathscr{C}^{\infty}(C)$ denote the commutative subalgebra generated by $\{I_{0},\dots,I_{n}\}$. We refer to $\mathcal{A}_{\mathrm{cl}}$ as the Jacobi commutative algebra of the integrable system. This terminology reflects the fact that $\mathcal{A}_{\mathrm{cl}}$ plays, for contact dynamics, the same role as the Poisson-commutative algebra of integrals in Liouville-integrable Hamiltonian systems: it encodes the distinguished observables whose mutual relations are governed by the underlying Jacobi structure.
@@
A contact system is naturally encoded by its symplectization (see Propoisition 2.8). In particular, the contact Hamiltonian dynamics on $C$ is the reduction (or, equivalently, the restriction to a transverse section $r=\mathrm{const}$) of the homogeneous Hamiltonian dynamics on $M$.
@@
For this reason, the correct correspondence between classical and quantum structures should be understood as follows: the quantization map $Q_{\hbar}$ assigns operators to the nonconstant generators of the Jacobi-commutative algebra $\mathcal{A}_{\mathrm{cl}}$, while the unit is added automatically on the quantum side. The non-trivial content of contact compatibility and of quantum constants of motion therefore lies entirely in the quantization of the dissipated quantities, not in the unit element.
@@
Let $\mathcal{A}_{\mathrm{cl}}=\mathrm{span}\{I\}$ be the Jacobi-commutative subalgebra generated by a single dissipated quantity $I$. Any semiclassical quantization $Q_{\hbar}$ maps the non-trivial generator $I$ to a self-adjoint operator $\widehat{I}_{\hbar}$. The corresponding quantum algebra is then taken to be the unital $C^{\ast}$-algebra
@@
A contact-compatible Lindblad generator is precisely one whose Heisenberg adjoint leaves $\mathcal{A}_{\hbar}$ pointwise invariant and satisfies $\mathcal{L}^{\dagger}(\widehat{I}_{\hbar})=0$. In this minimal setting, contact compatibility reduces to the existence of a non-trivial invariant commutative subalgebra generated by a single dissipated quantity beyond the scalars. This example highlights a general structural feature: while the constant function $1$ plays a non-trivial role in the Jacobi algebra at the classical level, its quantum counterpart is automatically invariant. The content of contact compatibility therefore lies entirely in the quantization of the non-constant dissipated quantities.
@@
$\{I_{i},I_{j}\}_{\alpha}=0\quad\text{for all }i,j,\qquad\{h,I_{i}\}_{\alpha}=0$
for some contact Hamiltonian $h$.
@@
Therefore, the operators $\widehat{I}_{i,\hbar}$ generate a commutative $C^{\ast}$-subalgebra $\mathcal{A}_{\hbar}\subset\mathcal{B}(\mathcal{H}_{\hbar})$, which is unital by construction, since any $C^{\ast}$-subalgebra of $\mathcal{B}(\mathcal{H}_{\hbar})$ contains the identity operator.
@@ 4.4 Bi-Lindblad structures
Let $(C,\alpha,h;I_{0},\dots,I_{n})$ and $(\mathcal{H}_{\hbar},Q_{\hbar})$ be as in Definition 4.14. Assume in addition that the contact system arises from a bi-Hamiltonian Poisson–Lie system $(G,\pi_{0},\pi_{1};H_{0},H_{1})$ as in Section 3, so that the classical vector field $X$ is Hamiltonian with respect to both $\pi_{0}$ and $\pi_{1}$, and the hierarchy of integrals $\{I_{j}\}$ ultimately comes from the common bi-Hamiltonian invariants on $G$.
@@
Heuristically, (BL1) mirrors the classical requirement that any linear combination $\pi_{\lambda}=\pi_{1}-\lambda\pi_{0}$ of compatible Poisson bivectors is still a Poisson tensor: here convex combinations of the two generators remain physically admissible Lindblad generators. Condition (BL2) captures the fact that the same family of classical integrals $I_{j}$ arises from the bi-Hamiltonian hierarchy and should therefore be conserved by both quantum evolutions. Finally, (BL3) encodes the requirement that, in the semiclassical limit, each branch $\mathcal{L}^{(a)}_{\hbar}$ recovers the contact dynamics associated with a different classical Hamiltonian $H_{a}$, in analogy with the way a bi-Hamiltonian vector field is Hamiltonian with respect to both $\pi_{0}$ and $\pi_{1}$.
@@
Verification of (CC1). Since the Heisenberg adjoint is linear,
@@
We illustrate Definition 4.21 on a two-level system $\mathcal{H}=\mathbb{C}^{2}$, where the invariant commutative sector is generated by $\sigma_{z}$. Fix parameters $\omega_{0},\omega_{1}\in\mathbb{R}$ and $\gamma>0$, and define two GKSL generators on $\mathcal{B}(\mathbb{C}^{2})$ (not forming a semiclassical family in $\hbar$, but sufficient to illustrate the algebraic content of Definition 4.21):
@@
Note that $\mathcal{L}^{(0)}$ describes a purely unitary evolution, while $\mathcal{L}^{(1)}$ corresponds to a unitary evolution plus pure dephasing in the $\sigma_{z}$ basis. For every $\lambda\in[0,1]$, the operator
@@
Thus a bi-Lindblad structure provides, for each $\lambda\in[0,1]$, a specific choice of contact-compatible Lindblad generator whose semiclassical limit reflects a different point on the classical bi-Hamiltonian pencil. In the Euler-top-type example of Section 5, we shall see explicitly how the quantum generators $\mathcal{L}^{(0)}_{\hbar}$ and $\mathcal{L}^{(1)}_{\hbar}$ can be chosen so that their common quantum integrals $\widehat{I}_{j,\hbar}$ quantize the contact dissipated quantities $I_{j}$ obtained in Section 3, thereby realizing a concrete bi-Lindblad structure in the above sense.
@@
Before stating the Egorov-type criterion, we briefly recall what we mean by semiclassical Weyl quantization in the standard pseudodifferential setting. Assume that the exact symplectic manifold $(M,\omega)$ is globally symplectomorphic to $\mathbb{R}^{2d}$ (with $d=n+1$) with canonical coordinates $(x,\xi)\in\mathbb{R}^{\mathrm{d}}\times\mathbb{R}^{d}$ and symplectic form ${\omega=\sum_{j=1}^{d}\mathrm{d}\xi_{j}\wedge\mathrm{d}x_{j}}$.
@@
in operator norm for $a,b\in S^{0}(M)$ under standard symbol assumptions. This is the pseudodifferential backbone behind the Dirac condition and the Egorov-type limit used below. We refer to [58, 18] for comprehensive accounts of semiclassical Weyl quantization, symbol classes $S^{m}$, the Moyal product, and the Egorov theorem. See also [23, 48] for the Weyl calculus and its microlocal properties. From the perspective of deformation quantization, the Moyal product and the Dirac condition are discussed for instance in [34].
@@
$(C,\alpha,h)=(\mathbb{R}^{3}(q,p,z),\,\mathrm{d}z-p\,\mathrm{d}q,\,h=z-p),\qquad I_{1}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=z.$
@@
$\mathcal{A}_{\mathrm{cl}}=\mathrm{Jacobi\text{-}alg}\langle I_{0}=h,\,I_{1}=z\rangle\subset\mathscr{C}^{\infty}(C).$
@@
As a consequence, the Lindblad operators belong to $\mathcal{A}_{\hbar}$, and hence
@@
$\widehat{I}_{0,\hbar}=\widehat{h}_{\hbar},\qquad\widehat{I}_{1,\hbar}=\operatorname{Op}_{\hbar}^{\mathrm{w}}(I_{1}^{\Sigma}),\qquad I_{1}=z.$